# Data-free Universal Adversarial Perturbation with Pseudo-semantic Prior

### v1

@@ Abstract
Data-free Universal Adversarial Perturbation (UAP) is an image-agnostic adversarial attack that deceives deep neural networks using a single perturbation generated solely from random noise, without any data priors. However, traditional data-free UAP methods often suffer from limited transferability due to the absence of semantic information in random noise. To address this, we propose a novel data-free universal attack approach that generates a pseudo-semantic prior recursively from the UAPs, enriching semantic contents within the data-free UAP framework. Our method is based on the observation that UAPs inherently contain latent semantic information, enabling the generated UAP to act as an alternative data prior, by capturing a diverse range of semantics through region sampling. We further introduce a sample reweighting technique to emphasize hard examples by focusing on samples that are less affected by the UAP. By leveraging the semantic information from the pseudo-semantic prior, we also incorporate input transformations, typically ineffective in data-free UAPs due to the lack of semantic content in random priors, to boost black-box transferability. Comprehensive experiments on ImageNet show that our method achieves state-of-the-art performance in average fooling rate by a substantial margin, significantly improves attack transferability across various CNN architectures compared to existing data-free UAP methods, and even surpasses data-dependent UAP methods.
@@ 1 Introduction
Figure 1: Diverse semantic contents in both a real-image and UAP: (a) A whole image from the ImageNet dataset (top) and our generated data-free UAP (bottom) from DenseNet-121, shown at iteration 900 during the training phase. The Top-1 class and its score are shown below each image. (b) Cropped regions from the whole image (top) and our UAP (bottom). Those regions contain diverse semantics that differ from the class of the original images.
@@
Inspired by this insight, we propose a novel data-free UAP method, called PSP-UAP, which generates pseudo-semantic priors from a UAP during training. To capture more diverse semantics in pseudo-semantic priors, we randomly crop and resize regions to extract semantic samples and treat them as images to fool. This approach effectively addresses the data-free constraint in UAP generation by leveraging richer semantic information inherent in the UAP, rather than relying solely on random noise. However, randomly cropped semantic samples often vary in informativeness, and simply averaging their activations can disrupt the generation of an optimal UAP. To mitigate this problem, we introduce a sample reweighting to focus on hard examples less deceived by an attack with a current UAP. To further improve the attack transferability, we incorporate input transformations [40, 43], primarily used in image-specific adversarial attacks, into our data-free UAP framework. This approach has remained unexplored in existing data-free UAP methods, as random priors lack semantic information, limiting its effectiveness. In contrast, our pseudo-semantic prior contains richer semantic content than random priors, enabling improved transferability to unknown models through input transformations.
@@
  • Our sample reweighting method prioritizes challenging examples during UAP training, minimizing the influence of uninformative samples that arise from random sampling within our pseudo-semantic prior.
@@ 3 Methodology
In this section, we present our motivation and approach for generating the pseudo-semantic prior. We then describe our sample reweighting strategy for optimizing UAP, followed by input transformations applied to semantic samples derived from the pseudo-semantic prior.
@@ 3.2 Pseudo-Semantic Prior
where ${z}$ and $\delta_{t}$ denote random noise sampled from $p_{{z}}$ and the UAP being trained in $t$-th iteration, respectively. The pseudo-semantic prior $p_{\mathbf{x}}$ denotes a set of adversarial examples with $\delta_{t}$ derived from the random prior distribution. $\mathcal{C}$ is a sampler that draws $N$ numbers of semantic samples ${x}_{n}$ from $p_{\mathbf{x}}$ by applying crop and resize operations. Specifically, we randomly crop a region of the UAP, and resize it to the original scale of the UAP size. We believe that leveraging the generated semantic information embedded in different regions of the UAP provides more effective guidance for successfully attacking target features than relying solely on random noise.
3.3 Sample Reweighting
Semantic samples are randomly drawn from the pseudo-semantic prior, leading to an imbalance in difficulty due to variations in semantic content; some samples are easily fooled by the UAP, while others are more difficult to deceive. To tackle this, we propose a new sample reweighting method that prioritizes harder-to-fool samples.
Specifically, we compute the weight of each sample using the KL divergence between the original input and its adversarial counterpart during training. We define the original distribution $P$ and the adversarial distribution $Q$ from each semantic sample and its corresponding adversarial example generated by the current UAP, respectively, as follows:
$P({x}_{n})=f({x}_{n}),$ (5)
$Q({x}_{n})=f({x}_{n}+\delta_{t}),$ (6)
where $f(\cdot)$ is the temperature-scaled softmax output of a surrogate model. Then, we calculate the weights of each semantic sample using the KL-divergence:
$w_{n}=D_{KL}(P({x}_{n})\|Q({x}_{n}))^{-1}.$ (7)
@@ 3.5 Overall Loss
We integrate the pseudo-semantic prior, sample reweighting, and input transformation into Eq. (2) to optimize the proposed PSP-UAP. The final loss function of our PSP-UAP is defined as follows:
@@ 4 Experiments
Our experiments are implemented on PyTorch with a single NVIDIA A6000 GPU. We set $\epsilon$ $=$ $10/255$ to restrict $\ell_{\infty}$-norm, the maximum iteration $T$ as $10,000$, and the saturation threshold $r$ to $0.001\%$, following the setting in TRM-UAP [21, 16]. For input transformations, we set $\theta$ $=$ $6$ for rotation, $\beta_{low}$ $=$ $0.8$ and $\beta_{high}$ $=$ $4$ for scaling, and $m$ $=$ $2$ for random shuffling. Moreover, the number of semantic samples $N$ is set to $10$. We define temperature parameter $\tau\in\{1.0,5.0,5.0,3.0,5.0\}$ corresponding to AlexNet, VGG16, VGG19, ResNet152, GoogleNet. For the additional CNN models, we set $\tau\in\{3.0,10.0,2.0,3.0\}$ corresponding to ResNet50, DenseNet121, MobileNet-v3-Large, Inception-v3. We follow the same curriculum learning and saturation check strategy of TRM-UAP. We set different ratios to use the activation of intermediate layers across different models. To ensure a fair evaluation, we find the optimal parameters for TRM-UAP on RN50, DN121, MB-v3-L, and Inc-v3 through our best efforts.
@@ 4.2 Evaluation on Black-Box Attack
We evaluate the transferability of our UAPs on commonly used CNN models in the black-box scenario. Table 2 represents the attack performance across different settings, with columns representing target models and rows indicating surrogate models used for crafting the UAPs. As shown in Table 2, PSP-UAP achieves superior results than other data-free attack methods across most models, with performance comparable to, but slightly below, TRM-UAP on GoogleNet. Notably, in strictly black-box settings (excluding the white-box scenario where GoogleNet attacks itself), PSP-UAP achieves an average $FR$ of 70.1% compared to TRM-UAP’s 69.6%, demonstrating better transferability.
@@
Figure 4: Visualization of UAPs crafted by PSP-UAP during training on DenseNet121, ResNet152, and AlexNet. From left to right, the UAPs are shown at iterations 100, 900, 1700, and the final UAP after training. Pixel values are scaled to [0, 255].
@@ 4.3 Ablation Study
Input Transformation to Other Data-free UAP.
We evaluate the effect of applying input transformations to semantic samples in our method against random noise within TRM-UAP [16], using the same transformation $T(\cdot)$. To ensure a fair evaluation, we set the number of samples $N=10$ for both random noise and semantic samples. Figure 6 shows that the $FR$ of TRM-UAP decreases when input transformations are applied. In contrast, applying transformations to semantic samples drawn from our pseudo-semantic prior increases the $FR$ compared to using semantic samples alone. We believe this improvement results from the richer semantic content in our pseudo-semantic prior compared to random noise, enabling the UAP to learn a wider variety of patterns through input transformations.
@@
Figure 6: Comparison of UAP performance with transformations applied to random noises in TRM-UAP [16] and to semantic samples drawn from pseudo-semantic priors in our method. $T$ denotes input transformation.
The Number of Semantic Samples.
We report the attack performance with respect to the number of samples $N$ in Figure 7 to analyze the impact of the sample size on our method. Empirically, we set $N=10$ as it achieves the highest attack performance, though results show generally stable performances across different sample sizes. Notably, even with $N=1$, our method achieves superior attack performance compared to other data-free methods.
@@
Figure 7: Ablation study on the hyperparameter $N$. The $FR$ (%) of UAPs generated with different numbers of semantic samples ($N$) is evaluated across various models.
@@ 5 Conclusion
In this paper, we proposed a novel data-free universal attack method, called PSP-UAP. To address the lack of prior knowledge in the target domain, we employed the UAP as a prior enriched with semantic information, allowing us to draw semantic samples directly from the pseudo-semantic prior. Additionally, we introduced sample reweighting to ensure a balanced attack across semantic samples. To further enhance the transferability of UAP, we applied the input transformation attack methods to the semantic samples. We demonstrated the exceptional transferability of our method by comparing PSP-UAP alongside both data-free and data-dependent universal attack approaches across various CNN models on the ImageNet validation dataset.
@@ Appendix A Additional Experiments on CNN Models
Impact of Each Proposed Component
@@
Impact of Epsilon
@@
We illustrate adversarial examples attacked by our generated UAPs using ResNet152 in Figure 9 with different $\epsilon\in\{8,10,16\}$. As expected, smaller $\epsilon$ values result in minimal degradation to the original image, whereas larger $\epsilon$ values highlight more artifacts introduced by the UAP. Similarly, as shown in Table 6, smaller $\epsilon$ values lead to lower performance compared to larger $\epsilon$ values.
We also visualize the final UAPs crafted for each model and the intermediate UAPs used during the training phase to generate the pseudo-semantic prior in Figure 11. As discussed in the main manuscript, visually diverse patterns can be observed across different iterations, even on the same surrogate model. This demonstrates that our method effectively crafts UAPs even in the absence of prior knowledge by generating diverse semantic samples.
@@ Appendix B Ablation Study on Hyperparameters
Ratio of Convolutional Layers
Temperature Parameters
After determining the optimal convolution layer ratio, we use it as a basis to find the temperature parameter $\tau$, used in Eq. (6) for the temperature-scaled softmax output, by incrementally increasing it from 1 to 10 in steps of 1. The results are shown in Figure 10, with blue lines representing the outcomes and the blue stars indicating the temperature values used in our experiments. Our observations indicate that variations in the temperature parameter $\tau$ have minimal impact on the results.
In Table 7 and Table 8, we report the performances of our PSP-UAP with a fixed temperature ($\tau=4$, referred to as PSP-I) alongside PSP-UAP with optimal temperature values (PSP-D) and TRM-UAP for comparison. Even with a fixed temperature, the performance difference is minimal, and our method consistently outperforms TRM-UAP by a significant margin. This highlights the robustness of our approach, achieving strong results over TRM-UAP even without tuning the temperature parameter.
Input Transformations
To evaluate the impact of hyperparameters for input transformation, we conduct an ablation study on hyperparameters for rotation, scaling, and shuffling. Specifically, we perform a comparative analysis on ResNet152 with $\theta\in\{3,6,9,12\}$, $\beta_{high}\in\{2,4,6,8\}$, and $m\in\{2,3,4,5\}$ to determine the optimal parameter values, extending the evaluation across other models. Each transformation experiment is conducted independently to isolate and analyze its impact on attack performance. To prevent excessive semantic information loss from extreme scaling, the minimal bound of the scale factor, $\beta_{low}$, is set to $0.8$. Based on the analysis in Figure 12, we set $\theta=6$, $\beta_{high}=4$, and $m=2$ in our experiments. Consequently, by integrating input transformation into a data-free framework, our method achieves high black-box attack transferability while preserving white-box performance as much as possible.
Appendix C Limitations and Discussions
@@
Figure 12: Hyperparameter analysis on input transformation techniques. Each transformation is individually applied to semantic samples during UAP training. Gray dashed lines represent the value used in our experiments.
Appendix D Reproducibility
Our code will be publicly released if our paper is accepted. For reproducibility, we have included the code in the supplementary materials, with detailed algorithms provided in the main manuscript and hyperparameters outlined in this supplementary document.

### v2

@@ Abstract
Data-free Universal Adversarial Perturbation (UAP) is an image-agnostic adversarial attack that deceives deep neural networks using a single perturbation generated solely from random noise without relying on data priors. However, traditional data-free UAP methods often suffer from limited transferability due to the absence of semantic content in random noise. To address this issue, we propose a novel data-free universal attack method that recursively extracts pseudo-semantic priors directly from the UAPs during training to enrich the semantic content within the data-free UAP framework. Our approach effectively leverages latent semantic information within UAPs via region sampling, enabling successful input transformations—typically ineffective in traditional data-free UAP methods due to the lack of semantic cues—and significantly enhancing black-box transferability. Furthermore, we introduce a sample reweighting technique to mitigate potential imbalances from random sampling and transformations, emphasizing hard examples less affected by the UAPs. Comprehensive experiments on ImageNet show that our method achieves state-of-the-art performance in average fooling rate by a substantial margin, notably improves attack transferability across various CNN architectures compared to existing data-free UAP methods, and even surpasses data-dependent UAP methods. Code is available at: https://github.com/ChnanChan/PSP-UAP.
@@ 1 Introduction
Figure 1: Diverse semantic contents in both a real-image and UAP: (a) Whole images from the ImageNet dataset (top) and our generated data-free UAP (bottom) using DenseNet-121, shown at iteration 900 during the training phase. The Top-1 class and its score are shown below each image. (b) Cropped regions from the whole image (top) and our UAP (bottom). Those regions contain diverse semantics that differ from the class of the original images.
@@
Inspired by this insight, we propose PSP-UAP, a novel data-free UAP method that generates pseudo-semantic priors from a UAP during training. To capture more diverse semantics in pseudo-semantic priors, we randomly crop and resize regions to extract semantic samples and treat them as images to be fooled. This approach effectively addresses the data-free constraint in UAP generation by leveraging richer semantic information inherent in the UAP, rather than relying solely on random noise. To improve attack transferability, we further incorporate input transformations [44, 47], commonly used in image-specific adversarial attacks, into our data-free UAP framework. This strategy has not been explored in existing data-free UAP methods, as random priors lack semantic information, limiting its effectiveness. In contrast, our pseudo-semantic prior contains richer semantic content, thereby enabling improved transferability to unknown models through input transformations. Moreover, since semantic samples obtained through random cropping and transformation vary in informativeness, we introduce a sample reweighting that prioritizes hard examples, which are less effectively deceived by the current UAP, to improve the overall effectiveness of the generated UAPs.
@@
  • Our sample reweighting prioritizes challenging examples during UAP training, by reducing the influence of uninformative samples produced by random sampling in our pseudo-semantic prior and input transformations.
@@ 3 Methodology
In this section, we present our motivation and approach for generating the pseudo-semantic prior. We then describe the input transformation applied to semantic samples derived from the pseudo-semantic prior, followed by our sample reweighting strategy for optimizing UAP.
@@ 3.2 Pseudo-Semantic Prior
where ${z}$ and $\delta_{t}$ denote random noise sampled from $p_{{z}}$ and the UAP being trained in $t$-th iteration, respectively. The pseudo-semantic prior $p_{\mathbf{x}}$ denotes a set of adversarial examples with $\delta_{t}$ derived from the random prior distribution. $\mathcal{C}$ is a sampler that draws $N$ numbers of semantic samples ${x}_{n}$ from $p_{\mathbf{x}}$ by applying crop and resize operations. Specifically, we randomly crop a region of $\mathbf{x}$, the sum of random noise $z$ and the UAP $\delta_{t}$, and resize it to the original scale of the UAP size. We believe that leveraging the generated semantic information embedded in different regions of the UAP provides more effective guidance for successfully attacking target features than relying solely on random noise.
@@ 3.3 Input Transformation
3.4 Sample Reweighting
Random processes involved in drawing semantic samples from the pseudo-semantic prior and applying input transformations lead to an imbalance in difficulty due to variations in semantic content; some samples are easily fooled by the UAP, while others are more difficult to deceive. To tackle this, we propose a new sample reweighting method that prioritizes harder-to-fool samples.
Specifically, we compute a weight for each sample using the KL-divergence between the transformed input and its adversarial counterpart during training. We define the original distribution $P$ and the adversarial distribution $Q$ for each semantic sample and its corresponding adversarial example generated by the current UAP $\delta$ as follows:
$P({x}_{n})=f(T({x}_{n})),$ (5)
$Q({x}_{n})=f(T({x}_{n}+\delta_{t})),$ (6)
where $f(\cdot)$ denotes the temperature-scaled softmax output of a surrogate model and and $T(\cdot)$ represents a randomly selected input transformation as described in Sec. 3.3. We then compute the weights for each semantic sample using the KL-divergence:
$w_{n}=D_{KL}(P({x}_{n})\|Q({x}_{n}))^{-1}.$ (7)
@@ 3.5 Overall Loss
We integrate the pseudo-semantic prior, input transformation, and sample reweighting into Eq. (2) to optimize the proposed PSP-UAP. The final loss function of our PSP-UAP is defined as follows:
@@ 4 Experiments
Our experiments are implemented on PyTorch with a single NVIDIA A6000 GPU. We set $\epsilon$ $=$ $10/255$ to restrict $\ell_{\infty}$-norm, the maximum iteration $T$ as $10,000$, and the saturation threshold $r$ to $0.001\%$, following the setting in TRM-UAP [23, 17]. For input transformations, we set $\theta$ $=$ $6$ for rotation, $\beta_{low}$ $=$ $0.8$ and $\beta_{high}$ $=$ $4$ for scaling, and $m$ $=$ $2$ for random shuffling. Moreover, the number of semantic samples $N$ is set to $10$. We define the temperature parameter $\tau\in\{1.0,5.0,5.0,3.0,5.0\}$ corresponding to AlexNet, VGG16, VGG19, ResNet152, GoogleNet. For the additional CNN models, we set $\tau\in\{3.0,10.0,2.0,3.0\}$ corresponding to ResNet50, DenseNet121, MobileNet-v3-Large, Inception-v3. We follow the same curriculum learning and saturation check strategy of TRM-UAP. We set different ratios to use the activations of intermediate layers across different models. To ensure a fair evaluation, we find the optimal parameters for TRM-UAP on RN50, DN121, MB-v3-L, and Inc-v3 through our best efforts.
@@ 4.2 Evaluation on Black-Box Attack
We evaluate the transferability of our UAPs on commonly used CNN models in the black-box scenario. Table 2 represents the attack performance across different settings, with columns representing target models and rows indicating surrogate models to craft the UAPs. As shown in Table 2, PSP-UAP achieves superior results than other data-free attack methods across most models, with performance comparable to, but slightly below, TRM-UAP on GoogleNet. Notably, in strictly black-box settings (excluding the white-box scenario where GoogleNet attacks itself), PSP-UAP achieves an average $FR$ of 70.1% compared to TRM-UAP’s 69.6%, demonstrating better transferability.
@@ 4.3 Ablation Study
Figure 5: Demonstrating that PSP serves as a universal strategy to other data-free UAP methods. Avgerage fooling rate (%) refers average FR (%) on AlexNet, VGG16, VGG19, ResNet152, and GoogleNet, with UAP crafted on ResNet152. P, T, and RW denote PSP, input transformation, and sample reweighting, respectively.
Impact of Input Transformation.
We evaluate the impact of applying input transformations to semantic samples in our method versus random noise on two data-free UAP methods (e.g., TRM-UAP [17] and GD-UAP [23]), to show that input transformations are effective when semantic information is present. For a fair evaluation, we set the number of samples $N=10$ for both random noise and semantic samples. As shown in Figure 5, applying input transformation directly to random noise leads to performance degradation, whereas integrating our pseudo-semantic prior (PSP) significantly boosts performance. Moreover, each baseline achieves its highest $FR$ when combined with our full approach, including sample reweighting. We believe that this improvement results from the richer semantic content provided by our PSP, which allows the UAP to learn a broader range of patterns through input transformation. This experiment also validates our approach as a universal strategy.
Figure 6: Ablation study on the hyperparameters for input transformation and the number of semantic samples. The hyperparameters used in our experiments are marked with gray dashed line.
Hyperparameter Analysis.
We conduct a hyperparameter analysis to evaluate how different settings for input transformation and the number of semantic samples affect our method’s attack performance on the ImageNet validation set (see Figure 6). With the exception of extreme scaling values (e.g., 8), the proposed method demonstrates stable performance across various hyperparameter values. Furthermore, performance remains consistent across different the number of semantic samples; notably, even with $N=1$, method outperforms other data-free approaches. Note that all hyperparameters used are chosen based on the ImageNet train set (See Appendix for further details).
@@
Figure 7: Visualization of UAPs crafted by PSP-UAP during training on DenseNet121 and ResNet152. From left to right, the UAPs are shown at iterations 100, 900, 1700, 2500, 3300, and the final UAP after training. Pixel values are scaled to [0, 255].
@@ 5 Conclusion
In this paper, we proposed a novel data-free universal attack method that leverages UAPs as a prior enriched with semantic information. This approach allows us to directly draw semantic samples from the pseudo-semantic prior, overcoming the lack of target domain knowledge. To further enhance transferability, we applied input transformation methods to these semantic samples. Additionally, we introduced sample reweighting to ensure a balanced attack across semantic samples. We demonstrated the exceptional transferability of our method by comparing PSP-UAP with both data-free and data-dependent universal attack approaches across various CNN models on the ImageNet validation dataset.
Acknowledgements.
This work was supported by the IITP grants (RS-2019-II191842 (3%), RS-2021-II212068 (2%), RS-2022-II220926 (50%)) funded by MSIT, and the GIST-MIT Research Collaboration grant (45%) funded by GIST, Korea.
@@ Appendix A Additional Experiments on CNN Models
Impact of Each Proposed Component.
@@
Robustness against Defenses.
In Table 7, we validate the robustness of our method against JPEG compression [7] and ensemble adversarially trained models, ens3-adv-Inc-v3 and ens-adv-Inc-Res-v2 [39]. Our method consistently shows higher robustness than TRM-UAP, with UAP crafted on ResNet152.
Attack JPEG compression Ensemble methods
 AN   VGG16   VGG19  RN152  GN Inc-v3ens3 IncRes-v2ens
TRM  53.57   58.38  53.86  39.94*  45.44 17.0 11.5
Ours  56.74   73.62  69.41  58.61*  62.00 19.8 13.1
Table 7: Robustness evaluation of our method and TRM against defense methods: JPEG compression and ensemble adversarially trained models.
Attack Model CN-B DeiT3 Others Model CN-B DeiT3 Others
TRM CN-B 40.71 10.10 34.67 DeiT3 14.18 6.73 36.01
Ours 86.57 13.94 59.49 19.98 9.54 43.05
Table 8: $FR$ (%) comparison for ConvNext-B (CN) and DeiT3. Others denotes the average $FR$ (%) on AlexNet, VGG16, VGG19, ResNet152, and GoogleNet.
Diverse Surrogate Models
We craft UAPs on various models in the main manuscript. To further demonstrate the effectiveness of our method on recent architectures, we additionally evaluate it using ConvNext-B [18] and DeiT3 [38] as surrogate models. As shown in Table 8, our method achieves strong performance on ConvNext-B but underperforms on DeiT3, which is consistent with the behavior observed in TRM. We attribute this to the fact that both our baseline and TRM were originally designed for CNN-based models, which may result in limited effectiveness on ViT-based architectures.
@@
We illustrate adversarial examples attacked by our generated UAPs using ResNet152 in Figure 9 with different $\epsilon\in\{8,10,16\}$. As expected, smaller $\epsilon$ values result in minimal degradation to the original image, whereas larger $\epsilon$ values highlight more artifacts introduced by the UAP. Similarly, as shown in Table 6, smaller $\epsilon$ values lead to lower performance compared to larger $\epsilon$ values. We also visualize the final UAPs crafted for each model and the intermediate UAPs used during the training phase to generate the pseudo-semantic prior in Figure 12. As discussed in the main manuscript, visually diverse patterns can be observed across different iterations, even on the same surrogate model. This demonstrates that our method effectively crafts UAPs even in the absence of prior knowledge by generating diverse semantic samples.
@@ Appendix B Ablation Study on Hyperparameters
Ratio of Convolutional Layers
Temperature Parameters
After determining the optimal convolution layer ratio, we use it as a basis to find the temperature parameter $\tau$, used in Eq. (6) for the temperature-scaled softmax output, by incrementally increasing it from 1 to 10 in steps of 1. The results are shown in Figure 10, with blue lines representing the outcomes and the blue stars indicating the temperature values used in our experiments. Our observations indicate that variations in the temperature parameter $\tau$ have minimal impact on the results. In Table 9 and Table 10, we report the performances of our PSP-UAP with a fixed temperature ($\tau=4$, referred to as PSP-I) alongside PSP-UAP with optimal temperature values (PSP-D) and TRM-UAP for comparison. Even with a fixed temperature, the performance difference is minimal, and our method consistently outperforms TRM-UAP by a significant margin. This highlights the robustness of our approach, achieving strong results over TRM-UAP even without tuning the temperature parameter.
Figure 11: Hyperparameter analysis on ImageNet train set for input transformation and the number of semantic samples. The hyperparameters used in our experiments are marked with gray dashed line.
Hyperparameter Search on ImageNet Train Set
We conduct experiments on a randomly selected subset of 1,000 images from the ImageNet train set to determine hyperparameters for input transformation and the number of semantic samples. As shown in Figure 11, our method demonstrates consistent performance across various transformation hyperparameters and exhibits a similar tendency to the results in Figure 6 of the main manuscript. In the case of the number of semantic samples, the performance is relatively low when $N=1$, which is likely due to sample imbalance caused by randomly sampling 1,000 images rather than using a dedicated validation set. Nevertheless, we conduct our main experiments using $N=10$, selected based on training set results and achieve the highest fooling rate on the validation set compared to other methods.
Appendix C Limitations and Discussions
Furthermore, while our method achieves strong performance on CNN architectures, it demonstrates limited attack transferability on ViT-based models. This limitation appears to be inherent to both data-free and data-dependent UAP approaches. As a direction for future work, we intend to explore black-box UAP strategies specifically tailored to ViT-based architectures.